OpenStudy (anonymous):

Which type of triangle has side lengths of 2 in, 5 in, and 6 in?

OpenStudy (anonymous):

what kind of answer you want? classified by its sides or by its angles?

OpenStudy (anonymous):

like iscosles ,acute

OpenStudy (anonymous):

classified by sides, this is scalene because no sides are of the same length. to determine if we have acute, obtuse, or right, you'll have to first determine if these 3 sides form a right triangle. you would essentially be using the converse of the pythagorean theorem... since 6 is the longest side, we'll let that be the "c" in "a^2 + b^2 = c^2"

OpenStudy (anonymous):

now, is 2^2 + 5^2 = 6^2 ??

OpenStudy (anonymous):

hello?